The kinetic energy of a mass is given by ½MV² of its speed is doubled by what factor does the energy change
1
Expert's answer
2019-10-17T11:14:49-0400

The kinetic energy


"K=\frac{Mv^2}{2}"

If the speed of mass is doubled, then

"K'=\frac{Mv'^2}{2}=\frac{M(2v)^2}{2}=4\times\frac{Mv^2}{2}=4K"

So, kinetic energy increases by factor 4.